OpenStudy (anonymous):

please someone help me out Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (s):

i have seen this one recently: -2(x+3)=-2(x+1)-4 but i am not sure who asked it before. it s with no solution

OpenStudy (s):

-2x-6=-2x-2-4 -2x+2x=-2-4+6 0x=-6+6 0x=0